determine whether the ordered pair is a solution of the given linear equation.
2x+5y=-8: (1, -2)

Answer by checkley77(12844)

answer is YES.
2x+5y=-8: (1, -2)
2*1+5*-2=-8
2-10=-8
-8=-8